OpenStudy (anonymous):

Find the intersection points of: x = sin(y) y^2 - pi^2 = x Hint please

OpenStudy (anonymous):

substitute equation 1 to equation 2 I think

OpenStudy (anonymous):

i would start with \[y^2-\pi=\sin(y)\]

OpenStudy (anonymous):

after that i would be stuck

OpenStudy (turingtest):

lol

OpenStudy (anonymous):

I just know the answer

OpenStudy (anonymous):

@satellite73 by the way he only ask for the hint xD

OpenStudy (anonymous):

Suppose I cannot use graphing calculator.

OpenStudy (turingtest):

I don't think any of us has given a good hint so far, because I don't think any of us knows how to solve it yet I would maybe start thinking about\[y=\sqrt{x+\pi^2}\]

OpenStudy (anonymous):

then i have no idea unless you want so start with \[\sin(x)-x^2+\pi=0\] and use newton ralphson

OpenStudy (anonymous):

y = arcsin(x), and y = \sqrt{pi^2 - x}?

OpenStudy (anonymous):

y^2 - pi^2 = sin(y)... y=pi :)

OpenStudy (anonymous):

OH!

OpenStudy (anonymous):

oooooooooooooooh it is \(\pi^2\) not \(\pi\)!!

OpenStudy (turingtest):

but that won't help that is clearly an answer by inspection.... if that counts I thought you wanted a method

OpenStudy (turingtest):

the "but that won't help" part was in reference to an earlier comment

OpenStudy (anonymous):

Yeah by inspection it would work. But the full question asks to find the area A of the region enclosed by the functions y = sin(y) and y^2 - pi^2 = x

OpenStudy (anonymous):

sorry, x = sin(y), not y = sin(y)

OpenStudy (anonymous):

then inspection is what you need !

OpenStudy (turingtest):

I guess :(

OpenStudy (anonymous):

Oh, then pi and -ve pi are the answers?

OpenStudy (anonymous):

I just realized that -pi would also give LHS = RHS

OpenStudy (turingtest):

yes

OpenStudy (anonymous):

Wait, so then the bounded region is -ve pi to pi, and the right function is sin(y) and the left function is y^2 - pi^2 = x

OpenStudy (turingtest):

yeah

OpenStudy (anonymous):

here is the picture, correct this time http://www.wolframalpha.com/input/?i=x^2-pi^2+%3D+sin%28x%29

OpenStudy (anonymous):

(i think, track record not so good this morning)

OpenStudy (anonymous):

Ok from there on, I got it. Thanks for your help. I do not know who to give medal too though.. :/

OpenStudy (anonymous):

not me!

OpenStudy (turingtest):

your link did not work for me sat http://www.wolframalpha.com/input/?i=plot%20x%3Dsiny%2C%20y%5E2-pi%5E2%3Dx&t=crmtb01